a stone is thrown vertically up with a velocity of 20m/s.
determine
A. the maximum height of the ball will reach to
B. the time of flight of the ball

Answer by Alan3354(69443)

a stone is thrown vertically up with a velocity of 20m/s.
determine
A. the maximum height of the ball will reach to
B. the time of flight of the ball
---------------
What ball? A stone was thrown??
-------------
You didn't give a function, so I'll use this one.
Assuming the ball/stone starts at a height of zero:
h(t) = -4.9t^2 + 20t + 0
A. the maximum height of the ball will reach
That's the vertex of the parabola, at t = -b/2a
t = -20/-9.8
t = 100/49 seconds, ~ 2.041 seconds
h(100/49) = -4.9*(100/49)^2 + 20*100/49
=~ 20.408 meters
------------------
B. the time of flight of the ball
It's 2x the time in A, same time falling as ascending.
= 200/49 seconds
=~ 4.082 seconds